OpenStudy (anonymous):

please help me with the answer i will give you a medal Parking in a student lot costs $1 for the first half hour and $1.25 for each hour thereafter. A partial hour is charged the same as a full hour . What is the longest time that a student can park in this lot for $8? The longest time is

OpenStudy (anonymous):

8-1=7 7=1.25x=5/4 x x=28/5=5.6=6 total hours =6+1/2=13/2 hours

OpenStudy (debbieg):

Let x be the number of hours AFTER the first hour. Then: \[1+1.25x=8\]\[1.25x=7\]\[x=5.6\] So the student can stay 6 hours (the whole part of the 5.6, plus the first hour). Notice that they can't stay a partial hour or they will be charged for a full hour. At 6 hours, they will pay: 1+1.25(5)=$7.25 An additional hour (or part of an hour) would put them over an $8 charge.

OpenStudy (anonymous):

okay so the answer is

OpenStudy (anonymous):

8

OpenStudy (debbieg):

It's in my post above... did you understand everything I said? What makes you say 8?

OpenStudy (anonymous):

yes i understood

OpenStudy (debbieg):

OK, the answer is there.... it isn't 8. :)

OpenStudy (anonymous):

no i was confused but i get it now okay thank you

OpenStudy (debbieg):

Sure, happy to help. :)

OpenStudy (anonymous):

complete hours are only 6

OpenStudy (debbieg):

Ooops, the $1 is for the first HALF hour. So the equation is the same, but x is the number of hours after the first HALF hour. Since x=5.6 then, you can park there for the first half hour, plus the 5 full hour part of 5.6, so a total of 5.5 hours. At 5.5 hours, they will pay: 1+1.25(5)=$7.25, and can't afford another full (or partial) hour.

OpenStudy (anonymous):

so the answer was wrong it is 5 hours

OpenStudy (debbieg):

It is 5.5 hours. The 5 full hours, plus the first 1/2 hour for $1. (I think the question is a bit confusing, in that it makes it sound like there are no "partial hours".... but in fact, you start with a half hour, THEN only full hours after that.)

OpenStudy (anonymous):

yes you are right